OpenStudy (anonymous):

help please? c; -(x+1)= -2 (5-x)

jimthompson5910 (jim_thompson5910):

I'll start you off -(x+1)= -2 (5-x) -1*x-1*1= -2*5-2*(-x) -x - 1 = -10 + 2x -x = -10 + 2x + 1 -x - 2x = -10 + 1 I'll let you finish up